# Stochastic Online Conformal Prediction with Semi-Bandit Feedback

### v1

@@ 4.1 Preliminary Results
Our first lemma shows that $G_{t}$ defined in (3) converges uniformly to $G_{t}^{*}$ defined in (2). Note that the randomness comes from the sequence of random variables $s_{j}=f(x_{j},y_{j}^{*})$.
Lemma 4.1.
@@
See Appendix A.1. ∎
@@
Lemma 4.2.
Suppose that $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$. Then, we have $\tau_{t}\leq\tau^{*}$ for all $t\in[T]$.
Proof.
See Appendix A.1. ∎
Our next lemma bounds the range of $\overline{G}_{t}(\tau_{t})$.
Lemma 4.3.
Suppose that $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$. Then, for all $t\in[T]$, we have
Proof.
See Appendix A.1
4.2 Main Result
Our main theorem shows that Algorithm 1 satisfies the desired properties, in particular, $\tilde{\mathcal{O}}(\sqrt{T})$ regret.
Theorem 4.4.
The expected cumulative regret of Algorithm 1 satisfies
$R_{T}\leq K\left(2\log T+4\sqrt{T\log T}+1\right)+4\phi_{max},$
Proof.
See Appendix A.1. ∎
(a) SQuAD
(b) ImageNet
Figure 1: Cumulative Regret. We show the cumulative regret of our algorithm on SQuAD and ImageNet. We fix the horizon at $T=3000$ and take the average of 10 runs. The shaded areas are 95% bootstrap confidence intervals of each algorithm, generated from 200 bootstrap runs. We set $\lambda_{1}=0.1$ and $\lambda_{2}=10$. The total number of exploration rounds of ETC and Con-ETC are chosen via a grid search. Greedy, ACI and ETC incur significantly more regret than our algorithm (SCS) on both datasets. Compared with Con-ETC, SCS achieves a marginal regret improvement that is indistinguishable from noise. Note that the curves of the greedy algorithm are convex in the figure; their empirical distributions $G_{t}$ become increasingly truncated, so the algorithm continues to greedily choose $\tau_{t}=\sup\{\tau\in\mathbb{R}\mid G_{t}(\tau)\leq 1-\alpha\}$. ACI also suffers the same problem as it cannot properly handle steps where the ground truth label is not observed.
5 Experiments
We evaluate our proposed approach on image classification and documents retrieval. All of the experiments are carried out on one Apple M1 machine.
5.1 Image Classification
First, we evaluate our algorithm’s performance on online image classification. Specifically, we apply the Vision Transformer model on the ImageNet dataset.
Vision Transformer (ViT) is a well-known computer vision model that serves as a popular baseline (Dosovitskiy et al., 2020). Given an image, ViT returns a pseudo-probability for each candidate label. We use ViT as our scoring function $f$ and define the score $s$ as the pseudo-probability assigned to the ground truth label.
We focus on the ImageNet dataset, a large crowd-sourced image classification dataset (Deng et al., 2009). [^1] The data was obtained from https://www.image-net.org/ with a custom and non-commercial license. We used the $16\times 16$ down sampled version. Each image from the dataset belongs to exactly one class out of the 1,000 candidate classes. Consequently, the cardinality of the label domain is 1,000 (i.e., $|\mathcal{Y}|=1000$). In the experiments, images arrive sequentially, and the learner aims to construct the smallest candidate label set that achieves the desired coverage. We finetuned ViT on a subset of ImageNet training data (720,000 images) to warm start the model.
5.2 Documents Retrieval
Next, we apply algorithm 1 to online documents retrieval. In particular, we use the popular Dense Passage Retriever (DPR) model on the well-known SQuAD question-answering dataset.
Dense Passage Retriever is a popular model for open-domain information retrieval Karpukhin et al. (2020). It leverages a dual-encoder architecture that maps questions and candidate documents to embedding vectors. Denote the space of questions and documents as $\mathcal{Q}$ and $\mathcal{D}$. DPR consists of a question encoder $E_{Q}:\mathcal{Q}\mapsto\mathbb{R}^{768}$ and a document encoder $E_{D}:\mathcal{D}\mapsto\mathbb{R}^{768}$. Given a question $q$ and a set of candidates documents $D\subseteq\mathcal{D}$, the likelihood of a document $d\in D$ being the ground truth document is captured by the similarity score $s_{q,d}$:
$s_{q,d}=\frac{\exp({E_{Q}(q)\cdot E_{D}(d)})}{\sum_{d^{\prime}\in D}\exp(E_{Q}(q)\cdot E_{D}(d^{\prime}))}.$
We use this DPR similarity score as our scoring function $f$. Then, our goal is to construct the smallest set of candidate documents while guaranteeing that they contain the ground truth document with high probability.
Our dataset is SQuAD question-answering dataset Rajpurkar et al. (2016), which is a popular benchmark for reading comprehension. Each question in SQuAD can be answered by finding the relevant information in a corresponding Wikipedia paragraph known as the context. The authors of DPR make a few additional changes to adapt SQuAD better for document retrieval. First, paragraphs are further split into multiple, disjoint text blocks of 100 words, serving as the basic retrieval unit (i.e., candidate documents). Second, each question is paired with ground truth documents and a set of irrelevant documents. [^2] In the data, all questions are paired with 50 irrelevant documents. We construct the candidate documents set by including all 50 irrelevant documents and 1 ground truth document. We exclude questions that have 0 ground truth documents. The data were obtained from DPR’s public Github Repo: https://github.com/facebookresearch/DPR with licenses CC BY-SA 4.0 and CC BY-NC 4.0 In our experiments, for each question, we include one ground truth document and all the irrelevant documents to create the set of candidate documents.
(a) Excess Prediction Set Size (ImageNet)
(b) Excess Prediction Set Size (SQuAD)
(c) Undercoverage Count (ImageNet)
(d) Undercoverage Count (SQuAD)
Figure 2: Undercoverage Count and Prediction Set Size. We show the undercoverage count and the excess prediction set size of our algorithm on ImageNet and SQuAD. We fix the horizon at $T=3000$ and take the average from 10 runs. The shaded areas are 95% bootstrap confidence intervals generated from 200 bootstrap runs. In Panels 2(a) and 2(b), the red dotted line represents the optimal set size, normalized to 0. Additionally, we take the rolling average of the excess prediction set size and truncate the $y$-axis in Panel 2(a) and Panel 2(b) for better interpretability. ACI, ETC, and greedy construct smaller prediction set sizes, but they fail to achieve desired coverage rate. While Con-ETC provides the same coverage guarantee as our algorithm, our excess prediction set size is 44% smaller on ImageNet and 45% smaller on SQuAD.
5.3 Baselines
We compare to four baseline strategies: greedy, explore-then-commit (ETC), conservative explore-then-commit (Con-ETC), and Adaptive Conformal Inference (ACI). The greedy strategy chooses
$\tau_{t}=\sup\{\tau\in\mathbb{R}\mid G_{t}(\tau)\leq 1-\alpha\}$
at every step, which cannot guarantee the $\alpha$ coverage rate, leading it to undercover significantly more than desired. The ETC strategy chooses $\tau_{t}=-\infty$ in the first $m$ steps, then commits to
$\tau_{t}=\sup\{\tau\in\mathbb{R}\mid G_{m}(\tau)\leq 1-\alpha\}.$
Next, we also consider a conservative explore-then-commit algorithm (Con-ETC) that guarantees $\tau_{t}\leq\tau^{*}$ with probability at least $1-2/T$. In contrast to ETC, Con-ETC commits to
after the exploration period—i.e., it commits to a conservative choice of $\tau$ that satisfies our coverage guarantee. The number of exploration rounds of ETC and Con-ETC are chosen via a grid search.
Finally, we compare our algorithm with Adaptive Conformal Inference (ACI) proposed by (Gibbs and Candes, 2021). The algorithm adjusts $\tau_{t}$ based on whether the ground truth label is in the previous round’s prediction set. The learning rate $\gamma$ is chosen from a grid search in a candidate set proposed in (Gibbs and Candès, 2022).
5.4 Metrics
We consider three metrics. First, we compute the cumulative regret for the following reward function $\phi$:
for some $0<\lambda_{1}<\lambda_{2}$. Note this loss imposes a larger penalty for undercovering compared to overcovering. Next, we count the number of times each algorithm chooses a cutoff greater than $\tau^{*}$:
Finally, we consider the average prediction set size. In particular, we consider
Excess Prediction Set $\text{Excess Prediction Set Size}=\lvert C_{t}\rvert-\lvert C_{t}^{*}\rvert,$
i.e., the difference between the constructed prediction set size and the optimal prediction set size. While our regret guarantees hold for $\phi$, in practice, we typically care about the excess prediction set size.
5.5 Results
We have implemented Algorithm 1 on ImageNet and SQuAD, and compared its performance with the four baselines. We choose $\alpha=0.9$ and $T=3000$ in all experiments. Figure 1 shows the cumulative regrets, and Figure 2 shows the prediction set sizes and the expected undercoverage counts. The red dotted lines in Figure 2(a) and Figure 2(b) represent the optimal set size, normalized to 0. We also include the 95% bootstrap confidence intervals that are generated from 200 bootstrap runs for all of the results.
On the image classification task, the cumulative regret accrued by our algorithm is 99.7% lower than greedy, 73.6% lower than ETC, 99% lower than ACI, and 0.05% lower than Con-ETC. The difference between Con-ETC and our algorithm is indistinguishable from noise as their confidence intervals overlap.
Next, greedy, ETC, and ACI construct smaller prediction sets than our algorithm but fail to maintain proper coverage rate. Greedy undercovers $2968$ times, ETC undercovers $1004$ times, and ACI undercovers $2635$ times. Although ACI is able to adjust the size of its prediction sets, the adjustment only improves the results marginally compared to greedy due to the truncated empirical distribution of the score. Lastly, while Con-ETC still achieves the desired coverage probability, our algorithm constructs prediction set size that is 44% smaller. One may notice that the curve of ACI in Panel 2(a) is lies above 00, (i.e., the prediction sets produced by ACI are larger than the optimal sets), but ACI still fails to maintain the desired coverage probability, as shown in Figure 2(c). We note that the curves in Panel 2(a) are rolling averages of prediction set sizes. Because ACI’s prediction set fluctuates between being too large and too small, it often undercovers. Nonetheless, its prediction set sizes are larger than the optimal sets on average.
On the documents retrieval task, our algorithm achieves again achieves $99.8\%$ lower regret than greedy, $74\%$ lower than ETC, $0.07\%$ lower than Con-ETC, and $99.9\%$ lower than ACI. The improvement over Con-ETC is indistinguishable from noise.
Similar to the previous results, greedy, ETC, and ACI achieve smaller prediction set sizes on average than our algorithm; however, they fail to achieve the desired coverage rate. Greedy undercovers 2968 times, ETC undercovers 1020 times, and ACI undercovers 2920 times. Lastly, Con-ETC achieves the same coverage probability as our algorithm, but our algorithm constructs prediction sets that are significantly smaller (i.e., 45% smaller) on average than Con-ETC.
These results demonstrate our algorithm’s effectiveness at constructing prediction sets for both image classification and document retrieval. Our algorithm consistently achieves the desired coverage rate and quickly converges to the optimal prediction sets. In comparison, the greedy, ACI, and ETC baselines fail to achieve the desired coverage rate, whereas the Con-ETC baseline constructs significantly larger prediction sets.
6 Conclusion
We have proposed a novel conformal prediction algorithm for constructing online prediction sets under stochastic semi-bandit feedback. We have shown our our algorithm can be applied to learn optimal prediction sets in image classification tasks and document retrieval tasks. Our experiments demonstrate that we maintain the desired $\alpha$ coverage level while achieving prediction set sizes that converge to the optimal sizes.
Appendix A Appendix
A.1 Proof of Theoretical Results
See 4.1
Proof.
Define the full empirical CDF of $s$ to be
$G^{\prime}_{t}(\tau)=\frac{1}{t}\sum_{j=1}^{t}\mathbbm{1}(s_{j}\leq\tau).$
In addition, define the following two events:
$A_{<}$ =$=\left\{\sup_{\tau<\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}$
$A_{\geq}$ =$=\left\{\sup_{\tau\geq\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\},$
First, consider $A_{\geq}$. For $\tau\geq\tau_{t}$, by definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}^{*}(\tau)=G^{*}(\tau)$ and $G_{t}(\tau)=G^{\prime}_{t}(\tau)$. Thus, by the DKW inequality, we have
$\mathbb{P}[A_{\geq}]$ $\leq 2e^{-2t\epsilon^{2}}.$
Next, consider $A_{<}$. By definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}(\tau)=G_{t}^{*}(\tau)=0$, so
$\mathbb{P}[A_{<}]=0.$
Thus, by a union bound, we have
$\mathbb{P}\left[\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right]$ $\leq\mathbb{P}[A_{\geq}]+\mathbb{P}[A_{<}]$
$\leq 2e^{-2t\epsilon^{2}},$
as claimed. ∎
See 4.2
@@ A.1 Proof of Theoretical Results
Thus, we have $\tau_{1-\alpha,k}\leq\tau^{*}$. Since $\tau_{k}\leq tau^{*}$, it follows that:
@@
See 4.3
@@
By Lemma 4.2, we have $\tau_{t}\leq\tau^{*}$, implying that $G^{*}(\tau_{t})\leq G^{*}(\tau^{*})$. For the upper bound, we have:
$\overline{G}_{t}(\tau_{t})$ $\leq G_{t}^{*}(\tau_{t})+2\epsilon_{t}$
$=G^{*}(\tau_{t})+2\epsilon_{t}$
$\leq G^{*}(\tau^{*})+2\epsilon_{t}$
$=1-\alpha+2\epsilon_{t}.$
@@
$\overline{G}_{t}(\tau_{t})$ $\geq\frac{\lfloor(1-\alpha)t\rfloor-1}{t}$
$\geq\frac{(1-\alpha)t}{t}-\frac{2}{t}.$
@@
$\overline{G}_{t}(\tau_{t})$ $\geq\frac{\lfloor(1-\alpha)t\rfloor}{t}$
$\geq\frac{(1-\alpha)t}{t}-\frac{2}{t},$
@@
By a union bound and by Lemma 4.1, we have
@@
$R_{T}$ $=\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}E\right]\cdot\mathbb{P}[E]$
$\qquad+\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}\neg E\right]\cdot\mathbb{P}[\neg E]$
$=\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\cdot\mathbb{P}[E]$
$\qquad+\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid\neg E\right]\cdot\mathbb{P}[\neg E]$
$\leq\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\cdot\mathbb{P}[E]$
$\qquad+\sum_{t=1}^{T}\mathbb{E}\left[2\phi_{max}\mid\neg E\right]\cdot\mathbb{P}[\neg E]$
@@
The first inequality follows from Assumption 2.2. Using Lemma 4.3, we can bound $X_{t}$ as follows:
@@
where the first inequality follows from Assumption 2.1 and the third inequality follows from Lemma 4.3. Thus, we have
$\sum_{t=1}^{T}X_{t}$ $\leq\sum_{t=1}^{T}\left\{\frac{2K}{t}+4K\cdot\sqrt{\frac{\log T}{t}}\right\}$
$\leq K\left[2\log T+1+4\sqrt{T\log T}\}\right].$

### v2

@@ 5 Conclusion
We begin by proving several helper lemmas, and then establish our main result. Our first lemma shows that $G_{t}$ defined in (3) converges uniformly to $G_{t}^{*}$ defined in (2). Note that the randomness comes from the sequence of random variables $s_{j}=f(x_{j},y_{j}^{*})$.
Lemma A.1.
@@
Let the full empirical CDF of $s$ be $G^{\prime}_{t}(\tau)=\frac{1}{t}\sum_{j=1}^{t}\mathbbm{1}(s_{j}\leq\tau)$, and define the events
$A_{<}$ =$=\left\{\sup_{\tau<\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}\qquad\text{and}\qquad A_{\geq}=\left\{\sup_{\tau\geq\tau_{t}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right\}.$
First, consider $A_{\geq}$. For $\tau\geq\tau_{t}$, by definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}^{*}(\tau)=G^{*}(\tau)$ and $G_{t}(\tau)=G^{\prime}_{t}(\tau)$. Thus, by the DKW inequality, we have $\mathbb{P}[A_{\geq}]\leq 2e^{-2t\epsilon^{2}}$. Next, consider $A_{<}$. By definition of $G_{t}^{*}$ and $G_{t}$, we have $G_{t}(\tau)=G_{t}^{*}(\tau)=0$, so $\mathbb{P}[A_{<}]=0$. Thus, by a union bound, we have
$\mathbb{P}\left[\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|>\epsilon\right]$ $\leq\mathbb{P}[A_{\geq}]+\mathbb{P}[A_{<}]\leq 2e^{-2t\epsilon^{2}},$
as claimed. ∎
@@
Lemma A.2.
Suppose $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$. Then, $\tau_{t}\leq\tau^{*}$ for all $t\in[T]$.
@@
Thus, we have $\tau_{1-\alpha,k}\leq\tau^{*}$. Since $\tau_{k}\leq\tau^{*}$, it follows that:
@@
Our next lemma bounds the range of $\overline{G}_{t}(\tau_{t})$.
Lemma A.3.
Suppose $\sup_{\tau\in\mathbb{R}}|G_{t}(\tau)-G_{t}^{*}(\tau)|\leq\epsilon_{t}$ for all $t\in[T]$. Then, for all $t\in[T]$, we have
@@
By Lemma A.2, $\tau_{t}\leq\tau^{*}$, implying that $G^{*}(\tau_{t})\leq G^{*}(\tau^{*})$. For the upper bound, we have
$\overline{G}_{t}(\tau_{t})\leq G_{t}^{*}(\tau_{t})+2\epsilon_{t}=G^{*}(\tau_{t})+2\epsilon_{t}\leq G^{*}(\tau^{*})+2\epsilon_{t}=1-\alpha+2\epsilon_{t}.$
@@
$\overline{G}_{t}(\tau_{t})$ $\geq\frac{\lfloor(1-\alpha)t\rfloor-1}{t}\geq\frac{(1-\alpha)t}{t}-\frac{2}{t}.$
@@
$\overline{G}_{t}(\tau_{t})$ $\geq\frac{\lfloor(1-\alpha)t\rfloor}{t}\geq\frac{(1-\alpha)t}{t}-\frac{2}{t},$
@@
By a union bound and by Lemma A.1, we have
@@
$R_{T}$ $=\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}E\right]\cdot\mathbb{P}[E]+\mathbb{E}\left[\sum_{t=1}^{T}|\phi(\tau^{*})-\phi(\tau_{t})|\biggm{|}\neg E\right]\cdot\mathbb{P}[\neg E]$
$=\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\cdot\mathbb{P}[E]+\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid\neg E\right]\cdot\mathbb{P}[\neg E]$
$\leq\sum_{t=1}^{T}\mathbb{E}\left[|\phi(\tau^{*})-\phi(\tau_{t})|\mid E\right]\cdot\mathbb{P}[E]+\sum_{t=1}^{T}\mathbb{E}\left[2\phi_{max}\mid\neg E\right]\cdot\mathbb{P}[\neg E]$
@@
The first inequality follows from Assumption 2.2. Using Lemma A.3, we can bound $X_{t}$ as follows:
@@
where the first inequality follows from Assumption 2.1 and the third inequality follows from Lemma A.3. Thus, we have
$\sum_{t=1}^{T}X_{t}$ $\leq\sum_{t=1}^{T}\left\{\frac{2K}{t}+4K\cdot\sqrt{\frac{\log T}{t}}\right\}\leq K\left[2\log T+1+4\sqrt{T\log T}\}\right].$
@@
Appendix B Additional Experiments
We consider two additional baselines. First, explore-then-commit (ETC) chooses $\tau_{t}=-\infty$ in the first $m$ steps, and then commits to
$\tau_{t}=\sup\{\tau\in\mathbb{R}\mid G_{m}(\tau)\leq 1-\alpha\}.$
Next, conservative ETC (Con-ETC) uses the same strategy, except it commits to
after the exploration period. In other words, it commits to a conservative choice of $\tau$ that satisfies our coverage guarantee, and also guarantees $\tau_{t}\leq\tau^{*}$ with probability at least $1-2/T$. The number of exploration rounds are chosen via a grid search. Results are shown in Figures 4, 5, & 6. As can be seen, ETC fails to achieve the desired coverage rate since it does not account for uncertainty; conversely, Con-ETC achieves very high regret since it does not adaptively choose $\tau_{t}$ over time.
(a) ImageNet
(b) SQuAD
(c) Auction
Figure 4: Cumulative Regret
(a) ImageNet
(b) SQuAD
(c) Auction
Figure 5: Coverage Rate
(a) ImageNet
(b) SQuAD
(c) Auction
Figure 6: Undercoverage Count